State the sample space for each other of the following random experiments.



a. tossing a coin



b. rolling a die



c. number of male children in a family of 2 children



d. identifying the blood type of an individual



e. determine the civil status of an individual

1
Expert's answer
2022-04-05T09:28:37-0400

"a:\varOmega =\left\{ Head,Tail \right\} \\b:\varOmega =\left\{ 1,2,3,4,5,6 \right\} \\c:\varOmega =\left\{ 0,1,2 \right\} \\d:\varOmega =\left\{ A,B,AB,O \right\} \\e:\varOmega =\left\{ Married,Single,Divorced,Widowed \right\} \\"